10ml of h2o2 solution react with ki solution to produce 0.5g of iodine. Calculate percentage purity....H2O2+2KI gives 2KOH+I2

Grade:12th pass

2 Answers

Arun
4 years ago
 
Write the balanced eqn. first,
H2O2   +   2 KI    ------------------>       I2   +  K2O   +    H2O
moles of I2 = 0.5/254 =  0.00196
from mole analysis,
moles of H2O2 would be: 0.00196
mass of H2O2 = 0.00196*34 = 0.066929 gm.
% purity =  0.066929/10 * 100 = 0.66929  %
Vikas TU
4 years ago